OpenStudy (anonymous):

Please help! Let theta be an angle in quad. 2 such that sin theta = 3/5. Find the exact values of sec theta and tan theta. Sec theta=? Tan theta=?

OpenStudy (anonymous):

First let us get the hypotenuse = √{(-5)²+3²} =√34 so cos Ө = -5/√34 since Ө is quadrant II angle sin Ө = 3/√34 sec Ө is reciprocal of cos Ө = √34/5 and tan Ө = -3/5

OpenStudy (amistre64):

the hypot is no sqrt34

OpenStudy (amistre64):

consider: sin^2 + cos^2 = 1

OpenStudy (anonymous):

don't forget the most famousest right triangle of the all the granddaddy right triangle, the mother of all right triangles

OpenStudy (amistre64):

a square?

OpenStudy (anonymous):

in fact it is so famous, that it leaves its imprint on any other right triangle with integer sides

OpenStudy (anonymous):



OpenStudy (amistre64):

all i can think of is: oh no, there goes tokyo, go go godzilla; which then leads me into , everybody was kungfu fighting

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Z6byK2Gbp5s

OpenStudy (amistre64):

mkty was on a good track except that sin is defined as over/hypot 3 is a leg and 5 is the hypot

OpenStudy (anonymous):

and therefore the other side is 4, as pictured now take your ratios from that one